OpenStudy (anonymous):

Please dont ingnore me. I just need the answer. Please. 8x+2 OVER 4 + 7x-2 OVER 10 = 3

OpenStudy (saifoo.khan):

That would be considered as cheating.

OpenStudy (anonymous):

I can't understand it. Atleast some explanation.

OpenStudy (saifoo.khan):

take LCm.

OpenStudy (saifoo.khan):

which will be 20.

jimthompson5910 (jim_thompson5910):

you can start by multiplying every fraction by the LCD 20 to get 5(8x+2) + 2(7x-2) = 3*20

OpenStudy (anonymous):

Finally, Thank you ..

OpenStudy (anonymous):

(8x+2)/(4)+(7x−2)/(10)=3 multiply all by 20 5(8x+2) + 2(7x-2)=60 40x+10+14x-4=60 54x=54 x=1